Suppose I have 40% of the 700,000 outstanding shares. How many board members can I elect if we are voting for 10 seats under cumulative voting?
Answer #1

Basic Understanding for Cumulative Voting

Cumulative voting is the procedure followed for electing a new director or board director. Normally, each shareholder is entitled to one vote per share multiplied by the number of directors. Shareholder can vote proportionally to the number of shares they hold and also split the vote across multiple candidate or apply them to just one candidate.

Given in question = 40% of 700,000 = 280,000 shares.

No of director/ seats = 10

Vote under cumulative voting = Number of shares held * No of seats

                                                          = 280,000*10

                                                     = 2,800,000 votes.

Above votes can be given to a single candidate or it can equally or unequally distribute among 10 candidates. Maximum candidate which you can vote for is 10 and minimum candidate which you can vote for is 1.

On numbers of member you can elect, please see below
Votes needed for needed =S*X/(D+1)
S=(outstanding shares)
X = no of directors you want to select)
D= (Total number of directors+1)
=(7000000*10)/(10+1)
=636,364 votes needed to elect one director.

Number of director you can elect = Total votes you have/ votes needed for one director
= 2800000*10/636364
= 4.40
In practical sense you can elect 4 members on board.